# RGBD Objects in the Wild: Scaling Real-World 3D Object Learning from RGB-D Videos

### v2

@@ 1 Introduction
In this paper, we propose to collect a new dataset that contains large-scale RGB-D object videos across diverse object categories and presented in the wild. Our dataset, namely WildRGB-D, covers 8500 tabletop objects across 44 categories in 20K videos. The videos are taken using iPhones to go around the objects in 360 degrees (see Figure 2 for visualization). Examples of the dataset are shown in Figure 1. There are three types of videos: (i) Single object video where there is only one object presented on the table; (ii) Multi-object video where there are multiple objects presented at the same time; and (iii) Hand-object video where there is a static human hand grasping the object. The collection of the WildRGB-D  dataset not only considers the cluttered background in the real world, but also the common scenarios where the objects are occluded by human hands.
@@ 2 Related Work
DTU [1] limited SfM mesh RGB N/A 124
CO3D [53] full SfM pcl RGB 50 19k
MVImgNet [82] limited SfM pcl RGB 238 220k
Objectron [2] limited SfM pcl RGB 9 15k
@@
OmniObject3D [73] limited scanner mesh RGB 190 6k
@@
Table 1: Comparison of WildRGB-D dataset with other 3D object dataset. Some datasets don’t provide video and we mark in “none”. Some only cover partial angles, which is marked in “limited”. Asterisk(*) means partial annotations. Depth Src. means where the depth information comes from, including CAD models, SfM COLMAP, scanner devices and depth sensor in iPhones. pcl is the abbreviation of point cloud.
@@
One representative kind of 3D object dataset is 3D synthetic dataset, like ShapeNet [5] and ModelNet40 [74], which consist of category-level objects. 3D-FUTURE [22] and ABO [14] datasets are typical of higher quality mesh with textures. [33] and [63] introduce real-world category-specific object datasets that mainly focus on birds and chairs respectively. DTU [1] and BlendedMVS [78] are datasets designed for multi-view stereo and lack category-level classification. Objectron [2] provides rich annotations but only partial videos are fully 360 degree covered. CO3D [53] is a large-scale category-level dataset that annotates camera poses and depths with SfM tool COLMAP [56], which doesn’t provide depth information in real-world scale. Pascal3D [75] contains real-world 3D objects with pose annotations and CAD models in limited categories. Datasets collected with specialized hardware (scanner, dome, etc.) like GSO [18] and OmniObject3D [73] have more accurate 3D geometry models. However, they don’t have RGBD wild object videos collected. In the aspect of RGBD object datasets, Wild6D [23] features RGBD image sequences and 6D pose annotations while lacking full 360 coverage and category types. Choi et al. [13] proposes RGBD object-centric datasets in 44 categories, but with limited camera annotations. As a comparison, our proposed WildRGB-D  dataset contains almost 8500 recorded objects and nearly 20000 RGBD videos recorded all 360 degrees in 46 common categories from well-known 2D datasets, all with real-world scale camera poses and object mask annotations as well as aggregated point clouds. We present the detailed comparison in Tab. 1.
@@
Neural Radiance Field (NeRF) [46] is a kind of scene representation based on MLPs. It takes in sampled points along each ray and outputs the density and color of each point, which are then aggregated by volume rendering to synthesize views. [3, 4, 47, 65] propose new changes to the original NeRF to improve the visual quality and [7, 21, 48, 59, 38] advance the NeRF efficiency. In order to generalize the NeRF representations to other scenes,  [81, 6, 69, 28, 42] learn latent 3D representations and priors from a bunch of existed scenes to help synthesize views across different scenes. Derived from original NeRF, [67, 79, 83, 15, 49, 72] leverage Sign Distance Function (SDF) and represent the 3D scene by implicit surface, which has a more clear object boundary definition. WildRGB-D  dataset comprises various category-level objects and scenes on a large scale, which is suitable for novel view synthesis benchmarks and helps boost more mature reconstruction algorithms and generalizable 3D scene representations.
@@ 3.1 Data Collections, Processing, and Annotation
In order to collect RGBD video on a large scale expediently and economically, we record with the help of iPhone front camera using Record3D  [^1] https://record3d.app/ App and rotate the camera around the object so that full 360-degree views of objects are captured with RGB images and the corresponding depth images. Camera rotating speed is controlled equably to ensure less blur in videos. We select 46 common categories from well-known 2D datasets [17, 36]. We record three videos for every selected object, which are composed of single-object video, multi-object video, and hand-object video. Every recorded video has been checked and some are left behind due to poor SLAM camera pose estimation. Details of WildRGB-D Dataset are listed in Fig. 3.
@@
Our WildRGB-D dataset has 3D annotations including camera poses in real-world scale, scene point clouds, and central object masks. In order to attain real-world scale camera poses, applying Structure-from-Motion (SfM) framework COLMAP [56] will result in relatively inaccurate SfM 3D camera pose annotations and the discrepancy of scaling in poses compared to real-world scaling, since SfM only leverages RGB image sequence and doesn’t take corresponding depths into account. The mature RGBD Simultaneous Localization and Mapping [19, 57] (SLAM) algorithms have the capability of exploiting the RGB images and depth information from depth sensor of mobile phones to reconstruct the 3D camera poses in real-world scale, which enables us to simply project the depth images and gain aggregated 3D point clouds (see Figure 5). Then we manually check the quality of the aggregated 3D point cloud and exclude videos in which SLAM fails to get accurate camera poses. To increase the probability of getting correct SLAM results for each video, we adopt two kinds of SLAM algorithms including BAD SLAM [57] and SLAM implementation from Open3D [87], which increase our successful rate to over 90%.
@@ 4.1 Novel View Synthesis
We select ten scenes from each category and uniformly sample images as validation split. We choose NeRF [46], Mip-NeRF 360 [4] and Instant-NGP [48] for evaluations. Results are shown in Tab. 2. We report the average PSNR, SSIM [71], LPIPS [86] and rendering depths Mean Average Error (MAE) compared with our sensor-collected depths across all categories. We also report metrics only related to the NVS quality of central objects using object masks. Results show that Mip-NeRF 360 and Instant-NGP outperform original NeRF in terms of visual quality metrics. NeRF-based methods perform better when we only concern with the recovery of central objects under object masks. What’s more, Mip-NeRF 360 performs best in learning single-scene geometry due to its special designs of proposal MLP and distortion loss. Visualization of single-scene NVS quality can be found in Figure 6. In brief, our dataset offers various and extensive categories and scenes for in-depth NVS experiments.
@@
Apart from single-scene NeRF optimizations, we also evaluate Generalizable NeRFs: Pixel-NeRF [81], MVSNeRF [6] and IBRNet [69] in the cross-scene setting. For each category in our dataset, we select the same ten test scenes as single-scene NVS experiments and train in the remaining scenes of the same category to learn per-category latent representations. We divide the 46 categories into three difficulty levels and report the average metrics of each level. For evaluation, we use three source views for every method to synthesize novel views. We report PSNR, SSIM and LPIPS to measure visual quality and depth MAE to measure the learned geometry quality. From Tab. 3, we observe that IBRNet outperforms in all three difficulty levels in terms of visual quality. Additionally, learned geometry quality isn’t highly correlated with the rendering visual quality in novel views. To sum up, our dataset provides great potential in learning category-level cross-scene NVS methods.
@@
Depth Supervised NVS
To benchmark our dataset in NVS tasks, we also study the influences that depth supervision brings. We choose Instant-NGP [48] in single-scene NVS methods and both Pixel-NeRF [81] and MVSNeRF [6] in cross-scene NVS methods. Since depth supervision in NeRF-based methods provides more information on the scene, it should be beneficial for them to learn better representations. Our experiment results in Tab. 4 prove that. In our experiment setting, we add L1 depth loss to every algorithm and choose the best-performance depth loss weight for them. Compared with conducting NVS tasks only with RGB images, the performances of both Instant-NGP and Pixel-NeRF get boosted when training with depth loss. As for experiments of MVSNeRF, since we’ve already added depth information in the original training as the guidance in building rays, it turns out that the improvements when training with extra added depth loss is limited. In a nutshell, with depth priors, not only does single-scene NVS get improved due to more knowledge of the scene learned in training, but cross-scene NVS methods also learn better generalization capabilities. Depth images in our dataset have the potential to boost NVS more accurate and generalizable.
4.2 Camera Pose Estimation
In this section, we benchmark two data-driven methods RelPose [84] and RelPose++ [34] for inference of the relative camera poses from multi-view images in a sparse setting. Leveraging the given annotations of camera poses and large-scale category-level video in our dataset, we aim to learn generalizable viewpoint inference capability from training-seen categories to unseen ones. Since WildRGB-D dataset has a full and dense 360-degree camera trajectory (visualized in Figure. 2), we can provide both a large-scale database and various view settings to assist training. In our experiments, we divided totally 46 categories into training and testing categories. We also hold some videos in training categories for evaluation. We adopt evaluation settings described in [84, 34] and report results in Tab. 5 and Tab. 6. We observe that these two methods can generalize well to other scenarios both in known categories and unseen categories since the relative rotation estimation errors are in a reasonable range (also see Figure 7 for visualization). However, the error of translation prediction is comparatively large in RelPose++, which still poses challenges in this field. To sum up, with extensive category-level object scenes in our dataset, WildRGB-D dataset can serve as large-scale training sources for generalizable camera pose estimation algorithms to achieve remarkable results.
4.3 RGBD Object Surface Reconstruction
In our experiment setting of object surface reconstruction, algorithms need to utilize RGBD image sequence and central object masks to reconstruct the surface mesh of the central object. Reconstruction only using RGB images is also evaluated for comparison. For evaluation of reconstruction quality, we calculate the Chamfer Distance between the reconstruction mesh and aggregated object point cloud which is derived from object-masked depth images. Ten single object scenes are selected in each category for evaluations of Instant-NGP [48] and Neusfacto [83]. From the results in Tab. 7, we observe that with depth priors, the quality of reconstruction is better. Additionally, the performance of Neusfacto with RGBD is superior to Instant-NGP. Visualization of Neusfacto RGBD reconstructions are shown in Figure 8. In brief, our dataset provides an RGBD object reconstruction evaluation track, which might give rise to the development of more mature algorithms in this field.
@@ 4.3 RGBD Object Surface Reconstruction
Figure 9: Object 6D pose estimation visualization. We visualize the predicted category-level 6D pose on three categories in Wild6D [23] test set (bottle, bowl and mug) using models that only perform self-supervised training on the corresponding category of WildRGB-D Dataset. The ground truth bounding boxes are colored in green, and the predicted bounding boxes are colored in red.
[…]

### v3

@@ 1 Introduction
In this paper, we propose to collect a new dataset that contains large-scale RGB-D object videos across diverse object categories and presented in the wild. Our dataset, namely WildRGB-D, covers 8500 tabletop objects across 44 categories in 20K videos. The videos are taken using iPhones to go around the objects in 360 degrees (see Figure 2 for visualization). Examples of the dataset are shown in Figure 1. There are three types of videos: (i) Single object video where there is only one object presented on the table; (ii) Multi-object video where there are multiple objects presented at the same time; and (iii) Hand-object video where there is a static human hand grasping the object. More video types add variety, creating occlusion for objects in scenes, which are worthy study cases in some tasks. The collection of the WildRGB-D  dataset not only considers the cluttered background in the real world, but also the common scenarios where the objects are occluded by human hands.
@@ 2 Related Work
DTU [1] limited COLMAP mesh RGB N/A 124
CO3D [54] full COLMAP pcl RGB 50 19k
MVImgNet [84] limited COLMAP pcl RGB 238 220k
Objectron [2] limited COLMAP pcl RGB 9 15k
@@
OmniObject3D [75] full scanner mesh RGB 190 6k
@@
Table 1: Comparison of WildRGB-D dataset with other 3D object dataset. Some datasets don’t provide video and we mark in “none”. Some only cover partial angles, which is marked in “limited”. Asterisk(*) means partial annotations. Depth Src. means where the depth information comes from, including CAD models, COLMAP, scanner devices and depth sensor in iPhones. pcl is the abbreviation of point cloud.
@@
One representative kind of 3D object dataset is the 3D synthetic dataset, like ShapeNet [5] and ModelNet40 [76], which consist of category-level objects. 3D-FUTURE [22] and ABO [14] datasets are typical of higher quality mesh with textures. [34] and [64] introduce real-world category-specific object datasets that mainly focus on birds and chairs respectively. DTU [1] and BlendedMVS [80] are datasets designed for multi-view stereo and lack category-level classification. Objectron [2] provides rich annotations but only partial videos are fully 360 degree covered. CO3D [54] is a large-scale category-level dataset that annotates camera poses and depths with COLMAP [57], which doesn’t provide depths in real-world scale, and MVImgNet [84] is also a dataset similar to CO3D. Pascal3D [77] contains real-world 3D objects with pose annotations and CAD models in limited categories. Datasets collected with specialized hardware (scanner, dome, etc.) like GSO [18] and OmniObject3D [75] have more accurate 3D geometry models and rendered depths from them. However, they don’t have RGBD wild object videos collected and lack real captured depths as well as background depths. In the aspect of RGBD object datasets, Wild6D [23] features RGBD image sequences and 6D pose annotations while lacking full 360 coverage and category types. Choi et al. [13] proposes RGBD object-centric datasets in 44 categories, but with limited camera annotations. As a comparison, our proposed WildRGB-D  dataset contains almost 8500 recorded objects and nearly 20000 RGBD videos recorded all 360 degrees in 46 common categories from well-known 2D datasets, all with real-world scale camera poses and object mask annotations as well as aggregated point clouds. We present the detailed comparison in Tab. 1.
@@
Neural Radiance Field (NeRF) [47] is a kind of scene representation based on MLPs. It takes in sampled points along each ray and outputs the density and color of each point, which are then aggregated by volume rendering to synthesize views. [3, 4, 48, 66] propose new changes to the original NeRF to improve the visual quality and [7, 21, 49, 60, 39] advance the NeRF efficiency. In order to generalize the NeRF representations to other scenes,  [83, 6, 70, 28, 43] learn latent 3D representations and priors from a bunch of existed scenes to help synthesize views across different scenes. Derived from original NeRF, [68, 81, 85, 15, 50, 74, 72] leverage Sign Distance Function (SDF) and represent the 3D scene by implicit surface, which has a more clear object boundary definition. Recently, 3D Gaussian Splating [31] has become a competitive alternative to NeRF. WildRGB-D  dataset comprises various category-level objects and scenes on a large scale, which is suitable for novel view synthesis benchmarks and helps boost more mature reconstruction algorithms and generalizable 3D scene representations.
@@ 3.1 Data Collections, Processing, and Annotation
In order to collect RGBD video on a large scale expediently and economically, we record with the help of an iPhone front camera using Record3D  [^1] https://record3d.app/ App and rotate the camera around the object so that full 360-degree views of objects are captured with RGB images and the corresponding depth images. Camera rotating speed is controlled equably by our collection setup to ensure less blur in videos. We select 46 common categories from well-known 2D datasets [17, 37]. We record three videos for every selected object, which are composed of single-object video, multi-object video, and hand-object video. Every recorded video has been checked and some are left behind due to poor SLAM camera pose estimation. Details of WildRGB-D Dataset are listed in Fig. 3.
@@
Our WildRGB-D dataset has 3D annotations including camera poses in real-world scale, scene point clouds, and central object masks. In order to attain real-world scale camera poses, instead of relying on COLMAP [57] to first generate camera poses and then the depths using the poses, we generate more accurate camera poses with the mature RGBD Simultaneous Localization and Mapping [19, 58] (SLAM) algorithm, which leverages our captured depths. Additionally, it has the capability of exploiting the RGB images and depth information from the depth sensor of mobile phones to reconstruct the 3D camera poses in real-world scale, which is different from COLMAP depths, which are not in real-world scale. It enables us to simply project the depth images and gain aggregated 3D point clouds (see Figure 5). Then we manually check the quality of the aggregated 3D point cloud and exclude videos in which SLAM fails to get accurate camera poses. To increase the probability of getting correct SLAM results for each video, we adopt two kinds of SLAM algorithms including BAD SLAM [58] and SLAM implementation from Open3D [89], which increase our successful rate to over 90%.
@@ 4.1 Novel View Synthesis
We select ten scenes from each category and uniformly sample images as validation split. We choose NeRF [47], Mip-NeRF 360 [4] and Instant-NGP [49] for evaluations. Results are shown in Tab. 2. We report the average PSNR, SSIM [73], LPIPS [88] and rendering depths Mean Average Error (MAE) compared with our sensor-collected depths across all categories. We also report metrics only related to the NVS quality of central objects using object masks. Results show that Mip-NeRF 360 and Instant-NGP outperform original NeRF in terms of visual quality metrics. NeRF-based methods perform better when we only concern with the recovery of central objects under object masks. What’s more, Mip-NeRF 360 performs best in learning single-scene geometry. Visualization can be found in Figure 6. In brief, our dataset offers extensive categories and scenes for in-depth NVS experiments.
@@
Apart from single-scene optimizations, we also evaluate Generalizable NeRFs: Pixel-NeRF [83], MVSNeRF [6] and IBRNet [70] in the cross-scene setting. For each category in our dataset, we select the same test scenes as single-scene NVS experiments and train in the remaining scenes of the same category to learn per-category latent representations. We divide the 46 categories into three difficulty levels and report the average metrics of each level. For evaluation, we use three source views to synthesize novel views. We report PSNR, SSIM and LPIPS to measure visual quality and depth MAE to measure the learned geometry quality. From Tab. 3, we observe that IBRNet outperforms in all three difficulty levels in terms of visual quality. Additionally, learned geometry quality isn’t highly correlated with the rendering visual quality in novel views. To sum up, our dataset provides great potential in learning category-level cross-scene NVS methods.
@@
Depth Supervised NVS
We also study the influences that depth supervision brings. We choose Instant-NGP [49] in single-scene NVS methods and both Pixel-NeRF [83] and MVSNeRF [6] in cross-scene NVS methods. Our experiment results in Tab. 4 prove that depth supervision is beneficial for these methods to learn better representations. In our experiment setting, we add L1 depth loss to every algorithm and choose the best-performance depth loss weight for them. Compared with conducting NVS tasks without depths, the performances of both Instant-NGP and Pixel-NeRF get boosted when training with depth loss. As for experiments of MVSNeRF, since we’ve already added depth information in the original training as the guidance in building rays, it turns out that the improvements when training with extra added depth loss are limited. In a nutshell, with depth priors, both single-scene NVS and cross-scene NVS methods learn better generalization capabilities, boosting NVS more accurate and generalizable.
4.2 Camera Pose Estimation
In this section, we benchmark two data-driven methods RelPose [86] and RelPose++ [35] for inference of the relative camera poses from multi-view images in a sparse setting. Leveraging the given annotations of camera poses and large-scale category-level video in our dataset, we aim to learn generalizable viewpoint inference capability from training-seen categories to unseen ones. Since WildRGB-D dataset has a full and dense 360-degree camera trajectory, we can provide both a large-scale database and various view settings to assist training. In our experiments, we divided totally 46 categories into training and testing categories. We also hold some videos in training categories for evaluation. We adopt evaluation settings described in [86, 35] and report results in Tab. 5 and Tab. 6. We observe that these two methods can generalize well to other scenarios both in known categories and unseen categories since the relative rotation estimation errors are in a reasonable range (also see Figure 7 for visualization). However, the error of translation prediction is comparatively large in RelPose++, which still poses challenges in this field. To sum up, WildRGB-D dataset can serve as large-scale training sources for generalizable camera pose estimation algorithms to achieve remarkable results.
4.3 RGBD Object Surface Reconstruction
In our experiment setting of object surface reconstruction, algorithms need to utilize RGBD image sequence and central object masks to reconstruct the surface mesh of the central object. Reconstruction without depths is also evaluated for comparison. For evaluation of reconstruction quality, we calculate the Chamfer Distance between the reconstruction mesh and aggregated object point cloud which is derived from object-masked depth images. Ten single object scenes are selected in each category for evaluations of Instant-NGP [49] and Neusfacto [85]. From the results in Tab. 7, we observe that reconstruction is better with depth priors. Additionally, the performance of Neusfacto with RGBD is superior to Instant-NGP, which shows that depths help the sdf-based method Neusfacto learn the correct object boundary and boost the performance more compared with Instant-NGP. The average deviation is high due to the varied reconstruction qualities across different categories in the dataset. Visualization of Neusfacto RGBD reconstructions are shown in Figure 8. In brief, our dataset provides an RGBD object reconstruction evaluation track, boosting the development of more mature algorithms in this field.
Figure 9: Object 6D pose estimation visualization. We visualize the predicted category-level 6D pose on three categories in Wild6D [23] test set (bottle, bowl, and mug) using models that only perform self-supervised training on the corresponding category of WildRGB-D Dataset. The ground truth bounding boxes are colored in green, and the predicted bounding boxes are in red.
@@ 4.3 RGBD Object Surface Reconstruction
Category Datasets IOU0.25 IOU0.5 5 deg. 5 deg. 10 deg. 10 deg.
@@ 5 Conclusion
Acknowledgment
This project was supported, in part, by the Amazon Research Award, the Qualcomm Innovation Fellowship, the Intel Rising Star Faculty Award, and the CISCO Faculty Award.